Suppose f(5)=7 and f(5.3)=7.9, approximate f'(5)
I can't seem to get started on this problem.
I am very grateful for all your help. Yours Faithfully, Michael.

Answer by stanbon(75887)

Suppose f(5)=7 and f(5.3)=7.9, approximate f'(5)
-----
Your have two points:: (5,7) and (5.3,7.9)
------
slope = (7.9-7)/(5.3-5) = 0.9/0.3 = 3
-----------------------
Ans: f'(5) = 3
----------------------
Cheers,
Stan H.
--------------